# MLLM-CL: Continual Learning for Multimodal Large Language Models

### v1

@@ C.3 Detailed Results of MR-LoRA
Table 35: Result matrix of MR-LoRA in domain continual learning
MR-LoRA RS Med AD Sci Fin
RS 80.82
Med 80.82 63.56
AD 80.74 63.01 52.49
Sci 80.80 63.22 52.49 53.55
Fin 79.88 62.69 51.84 52.45 89.69
Table 36: Result matrix of MR-LoRA in ability continual learning.
MR-LoRA OCR Math VP GA
OCR 35.00
Math 34.30 31.90
VP 35.80 32.20 62.06
GA 35.00 33.10 62.06 29.31
@@ D.1 Limitations
Although our study makes valuable contributions, we acknowledge the following limitations: (1) Model size and training limitations: This research focuses exclusively on MLLMs with 7 to 9 billion parameters. Owing to computational constraints, we did not explore larger models. (2) potential inaccuracies in StockQA dataset: Our StockQA dataset is generated by Qwen2.5-VL Qwen2.5-VL and the model may inadvertently produce inaccurate or misleading data. Moreover, biases inherent in the training data could manifest in the generated dataset, influencing the outcomes and interpretations of subsequent analyses. (3) Requirement for replay samples: Our MR-LoRA needs replay samples to train the MLLM enhanced selection router, a critical component responsible for optimizing decision-making during inference. While the number of samples we need is extremely small, it may be inconvenient in some special cases. For example, in cases where data privacy is a concern, the need for replay samples might cause problems. We hope to address these limitations in our future work to build a practical and lifelong-evolving MLLM.
@@ D.2 Broader Impacts
Appendix E Illustration of MLLM-CL Benchmark
@@
Figure 18: Examples of science task in domain continual learning.
@@
Figure 22: Examples of GUI agent task in ability continual learning.

### v2

@@ C.3 Detailed Results of MR-LoRA
Table 17: Result matrices of MR-LoRA in domain continual learning. LLaVA denotes LLaVA-based MR-LoRA, and InternVL denotes InternVL-based MR-LoRA.
LLaVA RS Med AD Sci Fin
RS 81.06
Med 81.06 65.73
AD 81.06 65.71 54.17
Sci 81.06 65.68 54.17 56.11
Fin 80.87 65.32 54.12 56.71 91.12
InternVL RS Med AD Sci Fin
RS 81.49
Med 81.49 66.40
AD 81.49 66.42 54.56
Sci 81.47 65.81 54.56 54.05
Fin 81.48 65.80 54.56 54.40 91.07
Table 18: Result matrices of MR-LoRA in ability continual learning. LLaVA denotes LLaVA-based MR-LoRA, and InternVL denotes InternVL-based MR-LoRA.
LLaVA OCR Math VP APP
OCR 33.60
Math 33.50 36.50
VP 33.50 36.40 64.97
APP 33.70 36.20 65.10 32.50
InternVL OCR Math VP APP
OCR 32.20
Math 33.80 36.40
VP 33.30 35.60 67.89
APP 33.00 35.70 67.51 33.75
@@ D.1 Limitations
Although our study makes valuable contributions, we acknowledge the following limitations: (1) Model size and training limitations: This research focuses exclusively on MLLMs with 7 billion parameters. Owing to computational constraints, we did not explore larger models. (2) potential inaccuracies in the StockQA dataset: Our StockQA dataset is generated by Qwen2.5-VL (Bai et al., 2025), and the model may inadvertently produce inaccurate or misleading data. Moreover, biases inherent in the training data could manifest in the generated dataset, influencing the outcomes and interpretations of subsequent analyses. We hope to address these limitations in our future work to build a practical and lifelong-evolving MLLM.
@@ D.2 Broader Impacts
Appendix E Inference Optimization with Caching
A key advantage of our method is its computational efficiency during inference. While our approach involves two distinct phases, we introduce a caching strategy that collapses the computational overhead. The most intensive operation—the forward pass through the backbone network (i.e., the visual encoder and LLM) is performed only once. We cache the resulting hidden states from each layer (specifically, the KV cache) after this single pass. Subsequently, our two lightweight modules, the router and the expert LoRA, operate sequentially on these cached states, obviating the need for a second full forward pass. This optimization reduces the computational cost from that of two full inferences to only marginally more than a single one, achieving a practical deployment cost comparable to standard single-pass methods, such as LoRA-FT (Hu et al., 2021).
Appendix F Use of LLM
In the preparation of this manuscript, we utilized a Large Language Model (LLM) in a capacity analogous to a conventional grammar-checking tool. Its application was strictly confined to copy-editing tasks, such as correcting spelling, improving grammar, and enhancing the clarity of author-generated text. No part of the research ideation, methodology, data analysis, or generation of substantive content was performed by the LLM.
Appendix G Visualization
G.1 Illustration of MLLM-CL Benchmark
@@
Figure 17: Examples of science task in domain continual learning.
@@
Figure 21: Examples of GUI agent task in ability continual learning.
@@
G.2 Visualization of Results
Fig.˜23 provides examples during DCL and ACL, respectively. We can find that some baselines like LoRA (Hu et al., 2021), MoELoRA (Chen et al., 2024a), HiDE (Guo et al., 2025a) overfit to the last learned task and output options that do not exist in domain continual learning. In ACL, most baselines, including HiDe (Guo et al., 2025a), DISCO (Guo et al., 2025b), CL-MoE (Huai et al., 2025), etc., miss part of their OCR ability and do not answer the question correctly.
Figure 23: Visualization of MR-LoRA and other baselines under domain continual learning and ability continual learning. The left part is testing the autonomous driving task after learning all domain tasks, while the right part is testing the OCR tasks after learning all ability tasks.